• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:04:49 on March21,2005

    Entry number 142115

    keyword=Day shift summary

    3/21/2005 day shift
    shift workers: X. Jiang, M. Potokar, J. P. Chen

    8:00 no beam, vaccum pump tripped
    8:35 beam back, resume data taking with 40 uA
    8:43 stopped beam,
    Beam study for next 6 hours
    9:15 Hall in controlled access, ED and company are working on the Septum.
    ? Hall back to beam permit
    13:25 ask for 5 uA to do a Harp Scan while right dipole still ramping up.
    13:30 did Harp scan run 2093.
    14:35 right dipole ready, asked for 40 uA, resume data taking on 12C.
    runs 2850-2852, wrong prescale factors.
    15:35 Folloing run coordinator's order, asked for beam to not going through Compton Chicane.
    16:00 still beam tuning.


    Production data on 12C
    run 2846, 7:13-8:07, 177 K
    run 2847, 8:35-8:43, 20K
    run 2848 cosmic
    run 2849, 10:27-13:15, cosmic run
    run 2850-2852 junk runs
    run 2853, 14:43-15:29, 200K.
    run 2854, 15:29-15:34, 1.3K.